You are here: Foswiki>Commissioning Web>Com20141118 (21 Nov 2014, TarasGolota)Edit Attach

TCS Build 2014D and other E-time tests - 20141118 UT

Observers: JHill, OKuhn, BRothberg (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support: (LBTO)
Telescope Support:
Software Support: KSummers, DSummers, CBiddick, MDeLaPena, TGolota, Igor (Tucson)
Instrument Support: JMorris (LBTO)
Engineering Support: DAshby

Summary

Sidereal guiding for LUCI1, LUCI2, MODS1, MODS2 works fine in this version of 2014D.

Non-sidereal guiding for LUCI2 works OK for rates of (XXXX).

Non-sidereal guiding for MODS1 was improved from previously, but still some jumps when the probe moves.

Non-sidereal guiding for MODS2 wouldn't guide in binocular mode, and wouldn't acquire the star in monocular mode.

Made some progress in understanding LBC focus.

Ran some WRS tests, but not much progress. It has problems with astigmatism and trefoil convergence.

Preparations

Details

LUCI1 monocular

00:42 ACQUIRE preset (sky a little bright) - IE=-39 CA=-41

00:45 ACTIVE preset - didn't find star (still too bright)

00:47 ACTIVE preset - guiding succeeds

00:52 changed guider_pixelsize to 0.0325 on LUCI_L.cfg and LUCI_R.cfg

LUCI1 and LUCI2 binocular

00:59 ACQUIRE preset binocular SX IE=-42 CA=-44 DX IE=-7 CA=-218

01:01 ACTIVE preset binocular - succeeds with lots of coma on each side

Collimation is running away on both sides. GCS is getting a "0" rotator angle.

David crashes PCS with a mis-pointed SYNC preset.

01:47 Repeating the bad preset did not crash PCS a second time.

Non-Sidereal LUCIs

01:52 Left side failed to move probe to -7,-7; Right side is running.

01:55 Send it again, but NS flag is cleared on right and set on left. This rejected the left preset, and hung on the right waiting for pointing corrections.

01:58 Michele cleared the preset waiting, but then the next preset still got stuck. Perhaps the previous rejected left preset was still there somewhere?

02:06 NS preset succeeds on asteroid Aethra (+5.8, -24.1) mag=14.89

Must use "irtc2" in the rdesktop command (The wiki instructions have now been updated.)

02:25 Non-sidereal preset Anahita succeeds but maybe guide probe is blocking IRTC (35, 9) mag=12.02

02:30 Restart GCSs to get Taras' patch of GCS that does the proper thing with the rotator angle.

02:32 ACTIVE sidereal preset to star in field of Anahita.

02:36 ACTIVE non-sidereal preset gs=1 for Anahita. Right side failed for GS 10e10 mm away???

02:39 ACTIVE non-sidereal preset gs=1 for Anahita - succeeds with asteroid on IRTC. It stayed within the 10 pixel box on IRTC for 5 minutes. (IRTC data was not saved.)

MODS1 monocular

02:55 Reconfigure swing arms both sides for MODS

03:10 ACQUIRE preset on left -- succeeds

03:11 ACTIVE preset on left --

Binocular MODS

03:xx ACQUIRE - first one failed because Olga homed the MODS2 stage while GCS is talking to it. Restarted GCS.

03:25 ACTIVE binocular preset

03:36 ACTIVE NS gs=1 to Aethra_0330 -- succeeds, but DX wfs seems permanently paused (with astigmatism present).

03:47:30 Reached Camera Limit on left side

              T.G.: Calibration tower was not retracted, moveProbe failed, correctly switched to Camera ROI mode (MODS_L does not allow it), activated StopGuiding, worked as expected.

03:50 monocular MODS

03:52 ACTIVE sidereal -- succeeds,

03:58 ACTIVE non-sidereal -- succeeds,

restart GCSR

04:37 ACTIVE non-sidereral Phaethon_20141118_0442 gs=1 (-103.5 -10.2) -succeeds with guiding jumps when the probe moves

04:45 changed wfs exposure time to 15000

04:48 prb_step_coeff 1.0 --> 1.8 (seems to make the star jump side-side rather than up-down ??)

04:52 prb_step_coeff 1.8 --> 1.0 (jumps switch back to mostly up-down)

05:10 ACTIVE non-sidereal Lacadiera_20141118_0518 gs=1 (-37.82, -16.25)

Binocular MODS

05:18 Authorize as MODS on both sides

05:19 ACQUIRE preset SX IE=-41 CA=-32 DX IE=+15 CA=-194

05:29 Cancel Sync Preset was needed to get our presets untangled

05:31 ACTIVE non-sidereal Lacadiera_20141118_0518 gs=1 (-37.82, -16.25) PCS is not getting any guide updates from the Right side.

MODS2 monocular

05:52 switch to MODS2 only authorization

05:56 ACTIVE sidereal on gs=1 of Lacadiera - succeeds

05:59 ACTIVE non-sidereal on gs=1 of Lacadiera - guide star is at the very bottom of the acq image - failed.

06:02 ACTIVE non-sidereal on gs=0 of Lacadiera - no star the acq image - failed.

06:04 ACTIVE sidereal on gs=0 of Lacadiera.....0600 - succeeds

06:08 ACTIVE non-sidereal on gs=2 of Lacadiera......0600 - no star the acq image - failed.

06:10 ACTIVE non-sidereal on gs=2 of Lacadiera......0600 - no star the acq image - failed.

                      T.G. See IT# 5393, incorrect AGw coefficients from MODS_R computer.:

LBC

06:15 Reconfigure to LBC on both sides

06:29 TRACK preset to WT10_297

Lots of testing of LBC version 2.1 here
From KS (19-Nov-2014):
The first test of LBC guiding ported to the CMU failed to find the FITS files. It was looking on /Repository/YYYMMDD/tec and they were not there in time. We decided the new archive software wasn't moving them from newdata as fast as the older system. So, Kellee added a mount to /newdata to the CMU and changed the code to look there.

The next test crashed in trackfirst in two separate attempts. We reverted to the operational software and continued testing.

Found the crashes were related to both red and blue running at the same time. The daytime testing never ran together. There was a software bug that caused by only a single instance of some structures by multiple threads.

07:45 back to "regular LBC software"

LBC Focus tests

08:15 dofpia on field at 03:18 49:41. Airmass 1.089. T=-2.3C. We need to move the hubs.

08:15 copoint

08:20 blue/red superfoc rV PA=0.
obsutil> !more bfoc0a.list
lbcb.20141118.082100.fits
lbcb.20141118.082149.fits
lbcb.20141118.082239.fits
lbcb.20141118.082328.fits
lbcb.20141118.082417.fits
lbcb.20141118.082507.fits
lbcb.20141118.082557.fits

08:30 blue/red superfoc rV PA=180, followed by PA=90 and then PA=270.

08:55 dofpia again on same field

08:55 superfoc at PA=180 and PA=0.

09:09 superfoc at PA=270 and PA=90. This will complete collection of data to assess the blue shim.

I examined the two PA=0 sequences and the first PA=180 sequence.
  • A plane should be fit to better measure the tilt, but it appears that at least nothing is worse. We still see the same bumps and dips in the focal plane.
  • I see some variation from dofpia to dofpia. Using fptilt1x.pl which measures the average/sig_n-1 of the stars within 500 pixels of the rotator center (a 3.75 arcmin box about the rotator center):
    • obsutil> !/home/lbto/kuhn/fptilt1x.pl SF_bfoc0a => SF_bfoc0a 0.0308 0.0049 40
    • obsutil> !/home/lbto/kuhn/fptilt1x.pl SF_bfoc180a => SF_bfoc180a 0.0051 0.0007 48
    • obsutil> !/home/lbto/kuhn/fptilt1x.pl SF_bfoc0b => SF_bfoc0b -0.0032 0.0005 42

Next up: measure focus offsets for dofpia which uses different regions of the chip 2
  • regular: 3201-4608
    • subject to curl at top of red chip 2
  • cent: 2201-3608
    • subject to bump between chips 2 and 3 of blue
  • low: 1201-2608
    • a new one to be tested tonight.

09:35 dofpialow to focus/collimate

09:38 RB_rV_sf7 (PA=0; bfoc0c.list).
/newdata/lbcb.20141118.093901.fits[2]   RB_rV_sf7       -0.200
/newdata/lbcb.20141118.093952.fits[2]   RB_rV_sf7       -0.140
/newdata/lbcb.20141118.094040.fits[2]   RB_rV_sf7       -0.080
/newdata/lbcb.20141118.094130.fits[2]   RB_rV_sf7       -0.020
/newdata/lbcb.20141118.094219.fits[2]   RB_rV_sf7       0.040
/newdata/lbcb.20141118.094310.fits[2]   RB_rV_sf7       0.100
/newdata/lbcb.20141118.094359.fits[2]   RB_rV_sf7       0.160

SF_bfoc0c -0.0027 0.0005 34

RB_rV_sf7 (PA=180; bfoc180c.list)
/newdata/lbcb.20141118.094637.fits[2]   RB_rV_sf7_pa180 -0.200
/newdata/lbcb.20141118.094726.fits[2]   RB_rV_sf7_pa180 -0.140
/newdata/lbcb.20141118.094816.fits[2]   RB_rV_sf7_pa180 -0.080
/newdata/lbcb.20141118.094905.fits[2]   RB_rV_sf7_pa180 -0.020
/newdata/lbcb.20141118.094955.fits[2]   RB_rV_sf7_pa180 0.040
/newdata/lbcb.20141118.095045.fits[2]   RB_rV_sf7_pa180 0.100
/newdata/lbcb.20141118.095134.fits[2]   RB_rV_sf7_pa180 0.160

SF_bfoc180c -0.0151 0.0029 29

RB_rV_sf7 (PA=0) same as before? (bfoc0d.list)
lbcb.20141118.095524.fits
lbcb.20141118.095614.fits
lbcb.20141118.095703.fits
lbcb.20141118.095754.fits
lbcb.20141118.095844.fits
lbcb.20141118.095933.fits
lbcb.20141118.100023.fits

SF_bfoc0d -0.0014 0.0003 26

10:15 dofpialow Some problems with dofpialow not loading the OB. John fixed a typo in the script. Running OK now.

10:20: rV PA=0 (bfoc0e.list). Will try next two other filter combinations: RUs and IB (added ig also) and the field at 05hr.
/newdata/lbcb.20141118.101753.fits[2]   RB_rV_sf7       -0.200
/newdata/lbcb.20141118.101847.fits[2]   RB_rV_sf7       -0.140
/newdata/lbcb.20141118.101937.fits[2]   RB_rV_sf7       -0.080
/newdata/lbcb.20141118.102029.fits[2]   RB_rV_sf7       -0.020
/newdata/lbcb.20141118.102119.fits[2]   RB_rV_sf7       0.040
/newdata/lbcb.20141118.102209.fits[2]   RB_rV_sf7       0.100
/newdata/lbcb.20141118.102259.fits[2]   RB_rV_sf7       0.160

SF_bfoc0e 0.0145 0.0029 26

10:25 RB_RUs_sf7

10:30 RB_IB_sf7

10:37 RB_ig_sf7

10:45 RB_rV_sf7 (again) SF_bfoc0f 0.0291 0.0063 22

10:53 slew to the 05hr field. dofpialow

11:00 rV
lbcb.20141118.110028.fits
lbcb.20141118.110116.fits
lbcb.20141118.110206.fits
lbcb.20141118.110256.fits
lbcb.20141118.110346.fits
lbcb.20141118.110435.fits
lbcb.20141118.110525.fits

SF_bfoc0g 0.0239 0.0098 7

11:15 IB

11:19 ig

11:25 rV

11:30 WRS testing: But the blue side stopped with an 'inhibit error'. Had to cycle software (lbckill/lbcstart). See the WRS wiki page for the testing log for this section.

LBC with DETXY OFFSET

12:56 Clear active optics on both sides

12:59 DOFPIALOW (from the lbcobs directory)

13:03 10 arcsec DETXY offset

13:07 90 sec exposures for guiding tests - only Blue took a 90 sec exposure. Were the guide exposures saturated? I'm not sure.

Close

13:12 Close for impending dawn

-- %USERSIG{JohnHill - 2014-11-18}%

-- %USERSIG{TarasGolota - 2014-11-20}%
Topic revision: r10 - 21 Nov 2014, TarasGolota
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback